Which expression is a proportion?
a) 2 = B
b) 2/5 = 6/x
c) 2/5 x 6/x
d) 12in./1ft.
Thanks for the help

Answer by AnlytcPhil(1807)

Which expression is a proportion? 
a) 2 = B

b) 2/5 = 6/x

c) 2/5 x 6/x

d) 12in./1ft.

Thanks for the help


A proportion is an equation with one fraction on each side, and
nothing else.

So the only one that is an equation with one fraction on each side
and nothing else, is choice (b).

The proportion is read

LEFT NUMERATOR is to LEFT DENOMINATOR as
 RIGHT NUMERATOR is to RIGHT DENOMINATOR

So the proportion

2/5 = 6/x

is not read

"Two fifths equals 6 over x" 

but instead as

"2 is to 5 as 6 is to x".

Edwin
AnlytcPhil@aol.com